OpenStudy (anonymous):

In the system shown below, what are the coordinates of the solution that lies in quadrant II? Write your answer in the form (a,b) without using spaces. x^2+4y^2=80 y= 1/32x^2

OpenStudy (apoorvk):

is that y = (1/32)x^2 ?? If yes then... from the second equation, x^2 = 32 y substitute this value for "x^2" in the first equation. You get: 32y^2 + 4y^2 = 80 Solve this^, get your (a,b) and the find out the quadrant!

OpenStudy (apoorvk):

I mean find out which root lies in the second quadrant. the 2nd quadrant has the abscissa(x-coordinate) negative and the ordinate(y-coordinate) positive.

OpenStudy (anonymous):

So when I solve for y in 32y^2 + 4y^2 = 80, I end up dividing 80 by 36 and then finding the square root of that, which ends up as a very messy number, correct? Which then turns the whole problem into a bit of a nightmare! Have I interpreted something incorrectly, or is this just a messy problem?

OpenStudy (apoorvk):

just a stupid messy problem, I guess. don't forget square roots will have have both the plus and minus values for 'y'. should get: y = +sqrt(20/9) and -sqrt(20/9) just confirming, which one is the second equation from below (am still doubting)